### v1

@@ 4.4.2 Layout-aware LiDAR Generation
To ascertain the layout-awareness of our LidarDM, we use CenterPoint [75] trained on real-world LiDAR scans to validate whether it can accurately detect objects from the LidarDM’s LiDAR scan. As LidarDM currently does not generate intensity, we trained CenterPoint excluding intensity of the real-world LiDAR scans, hence the potential mismatch between our results and those obtained in the original paper [75].
@@
Given an input layout $\mathcal{I}$, we generate the corresponding LiDAR scan, run CenterPoint on it, and evaluate using mean average precision (mAP) for vehicles. We compare this result with the mAP from the raw LiDAR. The result in Table 4 indicates that CenterPoint’s object detection on our generated LiDAR scan is comparable to the ground truth. We also compute the mAP agreement between our generated LiDAR scan and raw LiDAR scan, indicating a strong agreement between the two and demonstrating our approach’s map-awareness and realism.
4.4.3 Qualitative results
@@
We show qualitative results of our map-conditioned LiDAR sequence generation in Fig. 7. Our generated results closely match the map conditioning by adding flat surfaces and static vehicles that correspond to the provided layout. The use of physical-based LiDAR sensor simulation guarantees that the generated point clouds are properly occluded by obstacles and appear as a realistic LiDAR sweep pattern. We also showcase the accumulated LiDAR points over 90 frames in Waymax [20], highlighting LidarDM’s temporal consistency and map-awareness.
4.5 Augmenting Real Data with LidarDM
@@
LidarDM is the first LiDAR generative model capable of generating data conditioned on a given semantic layout. This capability offers the potential to augment the training data for 3D perception models, thereby further boosting their performance. To evaluate this capability, we first use LidarDM to generate around 70k frames of simulation data based on the layout from Waymo Dataset [63]. After that, we pre-train a LiDAR-based 3D object detection model, CenterPoint [75] (with PointPillars [36] as its backbone), on these generated LiDAR frames, paired with the object labels from the dataset. We then train the same model on 35k frames of real data, both with and without the pre-training stage on the simulation data, to test the benefits of the LidarDM-generated data. According to Table 4, LidarDM can act as an effective generative data augmentation strategy, offering more than a 3% improvement in detection accuracy.
5 Conclusion
@@
In this paper, we introduced LidarDM, a novel layout-conditioned latent diffusion model for generating realistic LiDAR point clouds. Our approach frames the problem as a joint 4D world creation and sensory data generation task and develops a novel latent diffusion model to create 3D scenes. The resulting point cloud videos are realistic, coherent, and layout-aware.
Limitations
So far, LidarDM relies on latent diffusion models, which are not real-time. Recent progress in latent consistency models promises to accelerate the generation process. We also leave LiDAR intensity modeling as future work.
5.0.1 Acknowledgement
This project is supported by the NCSA Faculty Fellowship, and NSF Awards #2331878, #2340254, and #2312102 and gifts from Intel, IBM, Amazon, and Meta. We greatly appreciate the NCSA for providing computing resources. Vlas Zyrianov is supported by the New Frontiers Fellowship. The authors thank Scott Lathrop and Aaron D. Saxton for technical support with the NCSA Delta cluster.

### v2

@@ IV-D2 Layout-aware LiDAR Generation
To ascertain the layout-awareness of our LidarDM, we use CenterPoint [62] trained on real-world LiDAR scans to validate whether it can accurately detect objects from the LidarDM’s LiDAR scan.
@@
Given an input layout $\mathcal{I}$, we generate the corresponding LiDAR scan, run CenterPoint on it, and evaluate using mean average precision (mAP) for vehicles. From our experiments, LidarDM achieves comparable mAP score (56.4%) compared to real LiDAR (59.7%), indicating strong semantic correlation of LidarDM’s point cloud to ground-truth. We compute the mAP agreement between our generated LiDAR scan and raw LiDAR scan to be up to 81.1%, showcasing a strong agreement between the two and demonstrating our approach’s map-awareness and realism.
IV-D3 Qualitative results
@@
We show qualitative results of our map-conditioned LiDAR sequence generation in Fig. 5. Our generated results closely match the map conditioning, and the accumulated points over 90 frames highlights LidarDM’s temporal consistency and map-awareness. The use of physical-based LiDAR sensor simulation guarantees that the generated point clouds are properly occluded by obstacles and appear as a realistic LiDAR sweep pattern.
IV-D4 Intensity Evaluation
To assess our intensity realism, we use Wasserstein distance (WD) to compare the intensity distribution of our generated point cloud with that of the real-world point cloud. As a baseline, we adopt the physics-based intensity model described in Sec. III-C3. Our experiment shows that LidarDM’s intensity distribution matches well with the gt (WD of 0.0197), which significantly outperforms the physics-based baseline (WD of 0.0813).
IV-E Augmenting Real Data with LidarDM
LidarDM is the first LiDAR generative model capable of generating data conditioned on semantic layouts. This capability offers the potential to augment the training data for and improve 3D perception and planning models.
IV-E1 Perception Model Sim2Real
@@
We first use LidarDM to generate around 70k frames of simulation data based on the layout from Waymo Dataset [54]. After that, we pre-train a LiDAR-based 3D object detection model, CenterPoint [62] (with PointPillars [63] as its backbone), on these generated LiDAR frames, paired with the object labels from the dataset. We then train the same model on 35k frames of real data, both with and without the pre-training stage on the simulation data, to test the benefits of the LidarDM-generated data. LidarDM-augmented model achieves mAP score of 61.3%, compared to 58.2% achieved with the real data-only model. This shows that LidarDM is an effective generative data augmentation strategy, offering over 3% improvement in detection accuracy.
IV-E2 Planning Model Sim2Real
Inspired by NMP [64], we developed a learning-based motion planner that takes the five most recent LiDAR observations (covering 0.5 seconds of past history) as input and generates 10 frame (with 0.3 second intervals between each) cost map of the car’s motion plan. We make two changes to the original NMP model: 1) our planner does not require privileged HD Maps as input, allowing the experimental results to focus on the quality of our generated LiDAR, and 2) the planner does not explicitly incorporate the ego car’s past trajectory [65, 66, 67]. We employ a soft cross-entropy loss to train the cost map and sample from a trajectory bank (generated from the Waymo dataset using K-Means) during inference.
To show LidarDM’s benefit for motion planning, we first train a model on 92k LidarDM-generated snippets, then fine-tune it on 9.2k real sequences. Expert driver trajectories serve as ground truth (GT) with traffic layout conditions generating LidarDM samples. For comparison, the same model is trained on only the 9.2k real sequences. Both models are trained for 30 epochs until convergence.
We report our results in Table III Using generative pre-training improves the performance of the planner in a low-data regime. In particular, our collision rate after 3 seconds has been reduced by $32\%$ (relative). To our knowledge, this is the first time conditional LiDAR generation has been shown to improve an end-to-end motion planner.
Config L2 (m) @ 1.0s L2 (m) @ 2.0s L2 (m) @ 3.0s Collision Rate (%)
9.2k Real $0.489$ $1.374$ $3.279$ $1.65\%$
9.2k Real+92k LidarDM $0.490$ $1.341$ $3.160$ $1.12\%$
TABLE III: Planner data augmentation: LidarDM-generated data can enhance performance of end-to-end planner. (  indicates best)
V Conclusion
@@
We presented LidarDM, a novel layout-conditioned latent diffusion model for generating realistic LiDAR point clouds. Our approach frames the problem as a joint 4D world creation and sensory data generation task and develops a novel latent diffusion model to create 3D scenes. The resulting point cloud videos are realistic, coherent, and layout-aware.